How many degrees are in the time 1:45 on a clock?
Answer by stanbon(75887)

How many degrees are in the time 1:45 on a clock?
---------------
Draw the picture
Let the position of the 12 correspond to zero degrees.
----------
The small hand is at (1/12)(360) = 30 degrees
The large hand is at (3/4)360 = 270 degrees
--------------
So the large angle formed by the hands is 270-30 = 240 degrees.
The small angle formed the hands is 30+90 = 120 degrees.
===========================================================
Cheers,
Stan H.